Without glasses, your Uncle Albert can see things clearly only if they are between 25 cm and 170 cm from his eyes. What power eyeglass lens will correct your uncle's myopia? Assume the lenses will sit 2.0 cm from his eyes. What is your uncle's near point when wearing these glasses?
Answer #1

object distance measured from retina

u = 170 - 2= 168

image distance, v = infinity

using lens formula

1/f = 1/ v - 1/u

1/f = 0 - 1/ u

1/f = - 1/ 168

f = - 168 cm

power, P = 100/f = - 100/168 = - 0.59 D

=========

b)

now image distance, v = - ( 25 - 2) = - 23 cm

using lens formula

1/f = 1/ v - 1/u

- 0.59 = - 1/0.23 +1 / u

u = 26.6 cm

========